What is the formula for power?


1
Expert's answer
2021-01-10T18:26:58-0500

Power (watt, W) "= \frac{Energy \;transfromed \; (joule, \;J)}{Time \;(seconds, \;s)}"

"P = \frac{E}{t}"

Power can also be calculated using the formula

Power (watt, W) = (Potential difference (volt, V)) "\times" Current (amp, A)

"P = V \times I"